* Test 32-channel pulsar modes
* Collect 32-channel RAW mode data
* Test fftshift values for 512-channel modes at L-Band
Ryan took detailed notes and then hit a bad button that caused the webpage to reset without saving. Lesson learned - save early and often. Blargh.
Less detailed summary:
The coherent 32 channel BOFs seem to have the wrong packet size (or at least unexpected packet size for guppi_daq) of 8208 instead 8224 bytes
The i1500x0032 mode seems OK
The levels and RFI seemed ok in the c0800x0512 and i0800x0512 modes. No major artifacts present, and no issue with diode SNR. In c0800x0512, banks C and D sometimes had the wrong number of subintegrations
There is a problem going from coherent pulsar --> 1 bank incoherent pulsar --> 4 bank spectral line mode. The one bank used in the incoherent mode seems to switch over to spectral line mode OK, but the banks that were previously in coherent mode seem to be partially stuck in coherent pulsar mode, i.e. the manager seems to think that it is in mode 1 but shared memory doesn't get reset to spectral line matrix hpc. Joe and Ray did some on the fly investigation but no conclusions yet.
Tried running some coherent mode scans while running the daq_server agent to monitor shared memory. In at least one scan, there were no issues, but in a few other scans one bank (Bank F once, and Bank C twice) didn't take data. No obvious configuration issues or errors in CLEO or Astrid.
Packet size issue in coherent 32 channel modes, but at least we can configure these modes now. Luke and Randy investigating.
Still having problems switching between spectral line and pulsar modes.